OpenStudy (anonymous):

A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks. Number of pages 50 100 150 200 Cost 1.75 2.55 3.35 4.15 Which equation gives the cost, y, in terms of the number of pages, x? A. y = 0.016x + 0.95 B. y = 0.016x + 1.75 C. y = 0.016x + 48.25 D. y = 52.6x + 0.95

OpenStudy (anonymous):

@chibiterasu

OpenStudy (anonymous):

@wendylisette

OpenStudy (anonymous):

y=0.016x+0.95 1.75=0.016(50)+0.95 1.75=1.75 2.55=0.016(100)+0.95 2.55=2.55 3.35=0.016(150)+0.95 3.35=3.35 4.15=0.016(200)+0.95 4.15=4.15